In a normally distributed data set of how long customers stay in your store, the mean is 44.8 minutes and the standard deviation is 2.6 minutes. Within what range would you expect 95% of your customers to stay in your store?

a2.6-44.8

b37.0-52.6

c42.2-47.4

d39.6-50.0
Answer #1

Here,

X ~ Normal(44.8, 2.62)

To find x_1 and x_2 such that

P(2.1 < X < .r2) = 0.95

) = 0.95 2,6 2.6 2.6

T1-44.8 P( KZ = 0.95 2.6

P(Z 22-44.8 2.6 )-P(Z-L 44.8 = 0.95

44.8 )-Ф( -2.0 T1-44.8 ) = 0.95

2 44.8 2.6 21-14.8 2.6 Ф( D( ) = 0.975-0.025

One solution may be

2 44.8 2.6 Ф( ) 0.975 and 44.8) 0,025 Ф(

Now,

2 44.8 2.6 Ф( ) 0.975

2 44.8 2.6 = Φ-1 (0.975) 1.95996

244.8(1.95996 2.6)

72 = 49.8959

and

44.8) 0,025 Ф(

Rightarrow rac{x_1-44.8}{2.6}= Phi^{-1}(0.025) = -1.95996

2.1 = 44.8-(1.95996 * 2.6)

1 39.7041

ans-> d) 39.6 - 50.0